# Optimizing Frequent Checkpointing via Low-Cost Differential for Distributed Training Systems

### v2

@@ V-C Optimizing Checkpointing Configuration
Therefore, we need to tune both FCF and BS to minimize LowDiff’s wasted time. For example, the wasted time in Table I achieves its minimum when $FCF=20$ and $BS=2$.
@@
The wasted time $T_{wasted}$ can be calculated by adding recovery overhead and steady state overhead (see §II-B):
$T_{wasted}$ $=\frac{NT}{M}(\frac{b}{2}+(R_{F}+\frac{R_{D}}{2}(\frac{1}{f\times b}-1)))+\frac{NT\times S\times f}{W}.$ (8)
To minimize $T_{wasted}$ on two variables $f$ and $b$, we take the first-order partial derivatives of $T_{wasted}$ with respect to $f$ and $b$.
$\begin{cases}\frac{\partial T_{wasted}}{\partial f}=\frac{NST}{W}-\frac{NT\times R_{D}}{2f^{2}\times M\times b}=0.\\ \frac{\partial T_{wasted}}{\partial b}=\frac{1}{2}-\frac{R_{D}}{2b^{2}\times f}=0.\end{cases}$ (9)
@@
The current design of LowDiff relies on gradient compression to enable efficient gradient reuse. However, in some high-precision training scenarios, any potential accuracy degradation caused by compression is unacceptable. If the distributed training system does not inherently employ a gradient compression strategy, making it impossible to leverage compression-induced bandwidth advantages. So we also need to design a non-compression high-frequency DC strategy for above scenarios. We next present two key insights on snapshotting and persistence, followed by a detailed description of how LowDiff can be enhanced in scenarios without gradient compression (LowDiff+ for short).
Figure 7: Layer-wise gradient reuse.
VI-A Layer-Wise Gradient Reusing and Snapshotting Strategy.
Issue of snapshotting. A key issue in the non-compression setting arises during the snapshot stage: unlike compressed gradients, transferring raw gradients without compression from GPU to CPU introduces heavier data movement and increases the likelihood of a stall. The snapshotting competes more directly with training for GPU and I/O resources, making overlap optimization even more critical to mitigate the training stall. To accelerate such large-scale snapshotting, we aim to design a fine-grained strategy that better overlaps gradient transmission with training.
@@
Insight 1: gradient can be reused layer-by-layer. We recall that gradients are produced incrementally in the backward pass, following the reverse layer order during training, as shown in Figure 7. Instead of waiting until the entire backward pass finishes to synchronize all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [deepspeed], PyTorch DDP [li2020pytorch], and Horovod [sergeev2018horovod]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced. Inspired by this execution model, we design a fine-grained reuse strategy: once a layer’s gradient has been synchronized, we immediately reuse it to the reusing queue, snapshot it to host memory, and stream it into a differential checkpoint assembler. This granular approach enables parallel data movement, allowing concurrent snapshot writes to CPU DRAM to better utilize memory bandwidth, thereby mitigating the high transfer volume. By pipelining backpropagation, communication, gradient reuse, and snapshotting at the layer granularity, we can effectively hide checkpointing overhead behind ongoing training, thereby maximizing overlap with training and reducing stalls.
Figure 8: Batching strategies of LowDiff and LowDiff+.
VI-B CPU-Based Asynchronous Persistence Strategy.
Issue of persistence. While the layer-wise strategy addresses the snapshotting issue in the non-compression setting, another issue remains: persistence of both full and differential checkpoints. The batching strategy described in §V-B accumulates small compressed updates on the CPU side before writing them out. It also periodically persists the full model state, which works effectively when update sizes are small and sparse. In the non-compression setting, differential checkpoints are significantly larger, and naively batching them demands a sustained write throughput that exceeds the available storage bandwidth, making it essential to design a more efficient checkpoint merging strategy that better lowers the writing data volume.
Insight 2: differential and full checkpoints can be batched together in CPU memory. We find that in the non-compression setting, the differential checkpoints (i.e. gradients) and full checkpoints do not need to be persisted separately. Instead, they can be directly fused with the full model state inside CPU memory, enabling the CPU to maintain an always-up-to-date replica of the GPU model. Once gradients are merged into the CPU-resident model state, persistence only needs to occur from the CPU model to storage. By fusing gradients into the CPU-resident model state instead of persisting them separately, redundant writes of differential checkpoints are avoided, which directly reduces the overall storage volume. The different batching strategies of LowDiff and LowDiff+ are shown in Figure 8. This strategy decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
@@ VI-C Design of LowDiff+
Design details. Based on Insight 1 and 2, we design LowDiff+ under non-compression scenarios. The architecture of LowDiff+ is shown in Figure 9, and its detailed algorithm is presented in Algorithm 2. In the training process, the model first performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer as shown in Figure 7 (Line 5). The layer-wise gradient communication is then handled by a communication thread pool $P_{g}$, where the gradient synchronization is assigned to multiple communication threads (Line 15). Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9). The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19). This parallel snapshotting allows different layer gradients to be written concurrently to CPU memory, thereby better utilizing the available DRAM bandwidth. We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state. After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12). While gradients are snapshotted layer-by-layer, the CPU-side update (Line 12) is performed once the full gradient is aggregated. This is because the Adam optimizer requires the complete gradient to correctly update its internal states (e.g., momentum and variance). The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [wang2023gemini]. Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead. Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients. The frequency of persistence is decided similarly to Checkfreq [mohan2021checkfreq], which ensures maximal overlapping between the persist and update of the model in CPU. This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
Figure 9: Architecture of LowDiff+.
Recovery of LowDiff+. When failure occurs, prior work categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures. For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one. In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time.
In contrast, software failures (e.g., software bugs) only cause the training process to terminate. This design assumes the failure is confined to the training processes (e.g., a CUDA error or NaN loss) and does not corrupt the separate checkpointing process’s memory, thus leaving the CPU memory state of the checkpointing process intact. In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints. This enables fast recovery under software failures, avoiding expensive checkpoint reloads and serving as a key optimization of LowDiff+.
Algorithm 2 Reusing in LowDiff+
1: GPU model state after the $t$-th iteration $M^{G}_{t}$, gradient of the $n$-th layer on node $i$ $G^{n}_{i,t}$, gradient communication and offload thread pools $P_{g}$ and $P_{s}$, reusing queue $Q$, corresponding handle sets $H_{g}$ and $H_{s}$.
@@
4:    for all $n$ do
5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$
6:      $H_{g}.append(P_{g}.execute(\textsc{Sync},G^{n}_{i,t}))$ 7:                                         $\triangleright$ Sync thread
8:    $G_{t}\leftarrow H_{g}.wait()$ $\triangleright$ Wait for sync
@@
11:   while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Zero-copy reuse
12:    $H_{s}.append(P_{s}.execute(\textsc{Snapshot},G^{n}_{t}))$ 13:                                       $\triangleright$ Snapshot thread
14:   $G_{t}\leftarrow H_{s}.wait()$ $\triangleright$ Wait for offload
15:   $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in CPU
16:   $C^{F}_{t}\leftarrow Save(M^{C}_{t})$ $\triangleright$ Persistence fully
17: function Sync Thread
18:   $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
19:   ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Zero-copy reuse
20:   return $G^{n}_{t}$
21: function Snapshot Thread
22:   $G^{n}_{t}\leftarrow Offload\_to\_cpu(G^{n}_{t})$ $\triangleright$ Snapshot
23:   return $G^{n}_{t}$
@@ VII-A Implementation of LowDiff
Reusing queue module. Since Python’s GIL prevents true multi-threading, LowDiff uses torch.multiprocessing with the spawn method for process isolation. The compressed gradient Reusing Queue is implemented via torch.multiprocessing.Queue(), which (1) preserves checkpoint order and (2) leverages CUDA IPC for zero-copy GPU memory sharing.
Batched gradient writing module. Differential checkpoints are offloaded from GPU to CPU memory with offload_to_cpu() (non-blocking) and temporarily cached in a CPU buffer. GPU memory handles are released after offloading. Once the batch size is reached, checkpoints are aggregated (tensor addition or dictionary accumulation) and persisted to local or remote storage.
@@
Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose the number of merge operations is $n$ with given $n$ differential checkpoints. Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints. Specifically, as shown in figure 10, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$). This parallelization reduces the number of merge operations between differential checkpoint and full checkpoint from $n$ to $\log n$. In Figure 10, the number of merges between differential checkpoint and full checkpoint is 2 (the serial process has $n=5$ merges). Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
@@ VII-B Implementation of LowDiff+
LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed]. Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning. For gradient communication, we leverage Python’s concurrent.futures library and employ a ThreadPoolExecutor to assign each layer-wise gradient to a communication thread, ensuring parallel transmission. The snapshotting thread pool is implemented in the same manner, handling the writing of gradients into CPU memory.
To maintain the CPU-side model replica, we configure PyTorch intra-op parallelism via torch.set_num_threads, setting the number of threads to match the available CPU cores in order to maximize parallel efficiency. Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation. Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [mohan2021checkfreq], thereby maximizing checkpointing throughput without interfering with GPU training.

### v3

@@ V-C Optimizing Checkpointing Configuration
Therefore, we need to tune both FCF and BS to minimize LowDiff’s wasted time. For example, the wasted time in Table I achieves its minimum wasted time when $FCF=20$ and $BS=2$.
@@
The wasted time $T_{w}$ can be calculated by adding recovery overhead and steady state overhead (see §II-B):
$T_{w}$ $=\frac{NT}{M}(\frac{b}{2}+(R_{F}+\frac{R_{D}}{2}(\frac{1}{f\times b}-1)))+\frac{NT\times S\times f}{W}.$ (8)
To minimize $T_{w}$ on two variables $f$ and $b$, we take the first-order partial derivatives of $T_{w}$ with respect to $f$ and $b$.
$\begin{cases}\frac{\partial T_{w}}{\partial f}=\frac{NST}{W}-\frac{NT\times R_{D}}{2f^{2}\times M\times b}=0.\\ \frac{\partial T_{w}}{\partial b}=\frac{1}{2}-\frac{R_{D}}{2b^{2}\times f}=0.\end{cases}$ (9)
@@ VI Enhancement
While LowDiff relies on gradient compression to achieve efficient gradient reuse, such compression may not be feasible or desirable in all training scenarios. In high-precision training tasks, gradient compression can introduce accuracy degradation, which is often unacceptable. Moreover, some distributed training systems may not employ gradient compression at all, thereby eliminating the bandwidth benefits that compression provides. To address these limitations, it is necessary to develop a high-frequency differential checkpointing (DC) strategy that does not depend on gradient compression.
In this section, we first outline two key insights regarding snapshotting and persistence mechanisms in the absence of compression. We then present LowDiff+, an enhanced version of LowDiff tailored for scenarios where gradient compression is unavailable or undesired.
(a) Snapshotting uncompressed gradient causes bandwidth contention and additional iteration time
(b) Persisting large volume batched gradients requires multiple iteration time
Figure 7: Issues of snapshotting and persistence.
VI-A Issues of Snapshotting and Persistence
Snapshotting issue. A key challenge in the non-compression setting arises during the snapshot phase: transferring uncompressed raw gradients from GPU to CPU introduces significantly heavier data movement compared to their compressed counterparts. This increased data traffic intensifies resource contention for GPU and I/O bandwidth, directly competing with training computation and substantially raising the risk of training stalls.
To quantify the impact of bandwidth contention, we measure the iteration time with and without snapshotting of uncompressed gradients, following the experimental setup in §III-A. As shown in Figure 7a, snapshotting uncompressed gradients increases the iteration time by 12.3%, 16.7%, 13.1%, and 17.4% for BERT-B, BERT-L, GPT-S, and GPT-L, respectively. In contrast, snapshotting compressed gradients incurs only a 2.4%–3.1% overhead. This substantial performance gap highlights the need for optimized overlap mechanisms to mitigate training stalls. To accelerate large-scale snapshotting without compression, we need to design a fine-grained, layer-wise gradient reuse strategy that better overlaps gradient transmission with ongoing training.
Figure 8: (Insight 1) Layer-wise gradient reuse.
Figure 9: (Insight 2) Batching strategies comparison.
Persistence issue. Although the layer-wise gradient reuse strategy mitigates the snapshotting overhead, a critical issue remains in the persistence phase when gradients are not compressed. The batched persistence strategy introduced in §V-B accumulates small compressed updates before writing them to storage, which is effective when updates are sparse and compact. However, without compression, differential checkpoints are substantially larger, and simply applying the same batching mechanism still demands write throughput that can easily saturate available storage bandwidth.
[…]
Figure 10: Architecture of LowDiff+.
VI-B Insights into Snapshotting and Persistence
@@
Insight 1: Gradient can be reused layer-by-layer. It is found that gradients are produced incrementally in the backward pass, following the reverse layer order during training [rumelhart1986learning]. Instead of waiting until the entire backward pass finishes to synchronize all gradients, which would incur high latency and risk blocking training, modern frameworks (e.g., DeepSpeed [deepspeed], PyTorch DDP [li2020pytorch], and Horovod [sergeev2018horovod]) already exploit layer-wise gradient availability by initiating communication as soon as each layer’s gradients are produced. Consequently, once a layer’s gradient has been synchronized, we can immediately reuse it in the reusing queue and snapshot it to host memory. This insight motivates a layer-wise fine-grain reuse and snapshotting design.
As shown in Figure 8, taking the 4-th layer as an example, once its local gradient $G^{4}_{i,t}$ is computed (Equation (2)) and synchronized (Equation (3)), the resulting global gradient $G_{t}^{4}$ is immediately enqueued for reuse and snapshotting. Crucially, this process forms a cascading pipeline: as the backward pass propagates to earlier layers (3, 2, and 1), their computation and synchronization overlap with the snapshotting of later layers. This fine-grained interleaving maximizes memory bandwidth utilization and effectively hides the checkpointing overhead within the training execution, thereby minimizing stalls.
Insight 2: Differential and full checkpoints can be merged incrementally. We observe that in LowDiff, differential checkpoints (i.e., raw gradients) are currently persisted separately from full checkpoints. Instead, they can be incrementally merged into the CPU-resident full checkpoint following the model update formula (Equation (4)), allowing the CPU to maintain an up-to-date replica of the GPU model. This merging strategy significantly reduces the volume of data to be persisted, as only the full model state, rather than individual differentials, needs to be persisted. This insight motivates an incremental merging design.
As illustrated in Figure 9, LowDiff persists both differential ($C^{B}_{t}$) and full ($C^{F}_{t}$) checkpoints into storage. In contrast, we can maintain a CPU-resident model replica ($M_{t}^{C}$). The model state in CPU ($M_{t}^{C}$) is initialized only once from the full model state in GPU ($M_{t}^{G}$) at the start of training. During subsequent iterations, incoming differential checkpoints ($C^{D}_{t}$) are immediately applied to update $M_{t}^{C}$ in memory, keeping it synchronized with the GPU model. Consequently, only the full model state in CPU ($C^{F}_{t}$) requires persistence. By merging the differentials into the CPU-resident state incrementally, this insight eliminates the overhead of writing separate differential checkpoints to storage.
@@ VI-C Design of LowDiff+
Algorithm 2 LowDiff+
1: GPU model state $M^{G}_{t}$ at the $t$-th iteration, gradient $G^{n}_{i,t}$ of the $n$-th layer on node $i$ , reusing queue $Q$, CPU model state $M^{C}_{t}$.
@@
4:    for all $n$ do $\triangleright$ Layer-by-Layer
5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$ $\triangleright$ Backward Pass
6:      $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
7:      ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Gradient Reuse
8:    $G_{t}\leftarrow\sum G^{n}_{t}$ $\triangleright$ Wait for Sync
@@
11:   for all $t$ do
12:    while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Gradient Reuse
13:      $C^{n}_{t}\leftarrow Snapshot(G^{n}_{t})$ $\triangleright$ Snapshot
14:    $G_{t}\leftarrow\sum C^{n}_{t}$ $\triangleright$ Wait for Snapshot
15:    $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Merge in CPU
16:    $C^{F}_{t}\leftarrow Persist(M^{C}_{t})$ $\triangleright$ Persist Regularly
Based on Insights 1 and 2, we design LowDiff+ under non-compression scenarios. The architecture of LowDiff+ is shown in Figure 10, and its detailed algorithm is presented in Algorithm 2.
Training Process. The model performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer (Line 4). Immediately following computation, each layer’s gradient undergoes synchronization (Line 5). Crucially, instead of blocking, the synchronized gradient $G^{n}_{t}$ is immediately enqueued into the reusing queue $Q$ (Line 6). This non-blocking operation allows the GPU to proceed with the backward pass of the next layer while the checkpointing process handles the data asynchronously. Finally, the training process aggregates the full gradient $G_{t}$ (Line 7) to update the GPU model state $M^{G}_{t+1}$ (Line 8).
Checkpointing Process. The checkpointing process is structured into two stages: layer-wise-reuse snapshotting (based on Insight 1) and incremental-merging persistence (based on Insight 2). For layer-wise snapshotting, the checkpointing process continuously fetches gradients $G^{n}_{t}$ from the reuse queue $Q$ (Line 11) and performs an immediate snapshot to host memory (Line 12). For incremental-merging persistence, we maintain a CPU-resident replica $M^{C}_{t}$. After all layers’ gradients are snapshotted, the full gradient is aggregated in the CPU memory (Line 13). Once aggregated, the reused gradients are applied to $M^{C}_{t}$ to update its model state (Line 14). The updated model state serves as the in-memory checkpoint. Since this update executes asynchronously alongside the next training iteration in the GPU, its latency is effectively hidden. Finally, persistence is performed asynchronously by writing the updated CPU model replica to storage (Line 15).
@@ VII-A Implementation of LowDiff
Reusing queue module. To bypass Python’s GIL, LowDiff uses torch.multiprocessing with the spawn method for process isolation. The Reusing Queue of compressed gradient is implemented via torch.multiprocessing.Queue(), which preserves checkpoint order and leverages CUDA IPC for zero-copy GPU memory sharing.
Batched gradient writing module. Differential checkpoints are offloaded from GPU to CPU memory with Snapshot() (non-blocking) and temporarily cached in a CPU buffer. GPU memory handles are released after offloading. Once the batch size is reached, checkpoints are aggregated (tensor addition or dictionary accumulation) and persisted to local or remote storage.
@@
Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose number of merge operations is $n$ with given $n$ differential checkpoints. Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints. Specifically, as shown in figure 11, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$). This parallelization reduces the number of merge operations between the differential checkpoint and the full checkpoint from $n$ to $\log n$. In Figure 11, the number of merges between the differential checkpoint and the full checkpoint is 2 (the serial process has $n=5$ merges). Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
@@ VII-B Implementation of LowDiff+
LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed]. The implementation of the three main modules of LowDiff+ is as follows:
Snapshotting module. To realize the layer-wise-reuse snapshotting execution, we leverage Python’s concurrent.futures library to manage concurrency through ThreadPoolExecutor. In the training process, we register backward hook for gradient synchronization, returning handles that allow the training loop to proceed non-blocking. On the checkpointing process, a separate thread pool concurrently retrieves gradients from the reusing queue and executes parallel snapshotting to the CPU DRAM. To prevent resource contention, we employ a priority-based policy that assigns a higher priority to gradient synchronization tasks over snapshotting tasks. This hierarchy ensures that in the event of bandwidth contention, the training task is always prioritized [zhang2025gockpt].
Persistence module. To maintain the CPU-resident model replica, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning. we configure PyTorch intra-op parallelism via torch.set_num_threads. Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation. Finally, the persistence to storage is performed asynchronously, following a design similar to Checkfreq [mohan2021checkfreq], which decouples disk I/O from the training critical path.
Recovery module. When failure occurs, prior work [gupta2024just, wang2023gemini] categorizes failures into two types: hardware failures and software failures. For hardware failures (e.g., GPU crashes) where the training state is lost, the distributed training system must replace nodes and reload checkpoints from persistent storage, incurring significant downtime. Conversely, software failures (e.g., NaN loss) only cause the training process to terminate but leave the host operating system intact. Our design isolates the checkpointing process from the training process. Consequently, a software failure in the training process does not corrupt the separate memory space of the checkpointing process. Recovery is achieved by reinitializing the training process and directly restoring the preserved, up-to-date CPU-resident model state to the GPUs. The recovery from software failures bypasses expensive disk reads, enabling rapid recovery for the most common failure modes.